Use the given confidence interval to find the margin of error and the sample proportion. (0.783,0.807)
Answer #1

Given that,
Lower bound= 0.783 ,

Upper bound= 0.807,

The sample proportion p^ can be calculated as,
p^ =( Upper bound +Lower bound)/2
=( 0.807+0.783)/2

= 1.59/ 2
=0.795

Margin of error = Upper bound - p^
= 0.807 - 0.795
= 0.012

Sample proportion= 0.795

Margin of error= 0.012